# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v2

@@ VI Future Directions
The emerging field of applying FM to FRS presents both opportunities and challenges. The combination of the powerful representation capabilities of FM and the data privacy protection advantages of FRS can enhance the performance of recommendation services and address some of the challenges that traditional FRS methods struggle to overcome. This section discusses some key future directions for FRS with FM, aiming to provide valuable guidance for future research and applications. These directions can offer higher quality personalized recommendation services while protecting user privacy. As technology continues to advance, it is reasonable to posit that FM will play an increasingly important role in future FRS.
@@ VI-A Data Augmentation
FMs, especially generative FMs such as LLMs and DMs, have the ability to generate training data needed for downstream tasks such as text and image generation [150, 151]. They can be used to generate more user interaction data, alleviating data scarcity problems. In FRS, data from different participants may be unbalanced or insufficient. Using FM for data augmentation can effectively increase the variety and amount of training data.
Through data augmentation, FM can help build more comprehensive user profiles and item characteristics, thereby improving the personalization and effectiveness of RS. For example, virtual interaction records of users or additional item descriptions can be generated to better capture user interests and preferences [152]. However, as mentioned earlier, the generated data may have quality issues that could degrade the performance of the RS and introduce biases and errors. Therefore, it is critical to perform rigorous validation and cleaning of the generated data to ensure that the synthetic data is statistically consistent with the real data and to reduce the noise introduced by the synthetic data [132]. Similarly, FM should be combined with privacy-preserving techniques, such as differential privacy, to ensure that no sensitive user information is revealed during the generation of synthetic data. In addition, FM should be used to generate diverse data to cover different scenarios and user behaviors, reducing the risk of bias and over-fitting [153].
@@ VI-B Cold-Start Recommendation
The core of the cold start problem is data sparsity, meaning that new users and new items lack sufficient historical interaction data [154]. This makes it difficult for the model to capture users’ true interests and preferences, which affects the accuracy and effectiveness of recommendations. Because FM has been pre-trained on large amounts of textual data, it has learned rich semantic information that helps generate high-quality representations of users and items. Therefore, FM has strong zero-shot and few-shot learning capabilities, allowing it to provide effective recommendations for cold start users and items even without sufficient interaction data [155, 41]. In addition, FM can transfer knowledge learned in one domain to another through techniques such as fine-tuning and prompt learning. This means that even in cold start situations, the model can use previously learned knowledge to make recommendations, thereby improving the performance of the recommendation system.
It is important to note that in a federated environment, collecting and using enough information to solve the cold start problem while protecting user privacy is a significant challenge. Ensuring the privacy and security of user data is a fundamental requirement for any FRS. In addition, multi-task learning [156] can improve the adaptability and generalization of basic models across different tasks and domains. For example, training on both recommendation and classification tasks simultaneously enables the model to make effective recommendations in more scenarios.
@@ VI-C Multi-Modal Recommendation
FM has the ability to process and understand different modalities of data, such as text, images, audio, and video [18]. In FRS, the integration of multi-modal data enables FM to build more comprehensive and multi-dimensional user profiles, thereby improving the personalization of recommendations [157, 158]. These rich user profiles can more accurately reflect the user’s interests and preferences, helping to deliver recommendations that better meet the user’s needs.
It is important to note that a major challenge of multi-modal data is its heterogeneity, meaning that different modalities of data have different characteristics and structures [159]. For example, text data is sequential, while image data consists of two-dimensional pixels. This heterogeneity increases the complexity of data fusion and processing. Therefore, the development of unified data representation methods that map different modalities of data into a common representation space is essential for model processing and integration.
In a federated environment, special attention must be paid to the privacy of multi-modal data. Multi-modal data may contain more sensitive information, such as the user’s voice or video, which requires more stringent privacy measures to effectively handle and protect this data. In addition, the processing and integration of multi-modal data requires significant computational resources and time. The training and inference processes of FM on multi-modal data can be more complex and time-consuming, making it critical to improve computational efficiency.
@@ VI-D Real-Time Recommendations
Real-time recommendation is a critical task in FRS, which aim to dynamically deliver personalized content based on real-time user behavior and contextual data [160]. This recommendation model emphasizes immediate processing and response to data, ensuring that users receive the most relevant recommendations in the shortest possible time. FM excels at understanding contextual information, which enhances their performance in processing user queries, item descriptions, and other textual data to improve recommendation accuracy and relevance.
However, real-time recommendations require processing and responding to user requests in extremely short time frames, which places high demands on computational resources and system latency. FM typically requires significant computational resources, which can lead to increased latency and negatively impact the user experience. In addition, limitations on the length of the input context can cause performance issues when dealing with long sequences of user behavior or a large number of candidate items.
In real-time recommendations, FM can also reflect biases from the training data in its recommendations, leading to unfair biases. For example, popular items may be preferentially recommended due to their high frequency in the training data, overshadowing long-tail items [136]. Therefore, it is crucial to explore efficient model compression and acceleration techniques, such as distillation [161] and pruning [162], to reduce computational complexity and latency. Introducing strategies such as sliding windows [163] can manage and optimize the input of contextual information, thereby improving the model’s performance in handling long behavioral sequences. Incremental learning methods [164, 165]can also be used to update the FM, and model distillation techniques can reduce model size, thereby lowering communication overhead.
VI-E Enhanced Recommendation Explainability
The federated nature and recommendation characteristics of FRS require transparent explanations of recommendation results to users. This transparency helps users understand the logic and reasoning behind the recommendations, thereby increasing their trust and satisfaction. Language FMs, pre-trained on large amounts of text data, have strong natural language generation capabilities. They can generate clear and coherent textual explanations while referencing relevant background knowledge, helping users to understand the reasons behind recommendations [166, 167]. For example, models such as ChatGPT can generate detailed recommendation explanations that explain why a particular item is suitable for the user [168]. However, generating detailed recommendation explanations often requires significant computational resources [169]. This is especially true for LLMs, which can increase latency and negatively impact the user experience. In addition, these LLMs may exhibit biases present in the pre-training data, resulting in unfair tendencies in the generated explanations, such as gender or race biases. Therefore, it is important to apply model compression and optimization techniques to reduce the computational burden of FM, and to design debiasing methods and fairness constraints to minimize biases in recommendation explanations. Furthermore, it is crucial to incorporate user feedback mechanisms to continuously improve the quality of recommendation explanations. By leveraging user feedback, strategies for generating recommendation explanations can be adapted and optimized, thereby increasing user satisfaction and trust.
VI-F Adavanced Metrics
Traditional evaluation metrics, which will be detailed in VII-B, are designed for specific tasks in recommendation scenarios, such as rating prediction or item ranking. However, the strength of FM in RS lies in its powerful generative capabilities, which allow it to generate items that have never appeared in historical data and recommend them to users. In addition, FM can generate explanations and user interaction content, which requires more complex and comprehensive evaluation criteria. In this context, how to evaluate the recommendation capabilities of these systems remains an open question. Therefore, FRS with FM requires new evaluation metrics tailored to the specific applications of FM in RS, especially in generative recommendation.
@@ VII-A Datasets
In FRS, the datasets used can be categorized into two types based on user feedback: explicit feedback and implicit feedback, as shown in Fig. 7. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data is clear and provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table I:
@@
  • Anime [^4] https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database: This dataset collects user preference data from the MyAnimeList website. It contains information from 73,516 users on 12,294 anime titles. Users can add anime to their completed list and rate them, and this dataset compiles these ratings.
@@
  • Douban [^6] https://www.kaggle.com/datasets/utmhikari/doubanmovieshortcomments: The Douban Movie dataset is a Chinese website where internet users can post their opinions and comments about films. This dataset contains over 2 million short comments on 28 movies from the Douban Movie website.
@@
  • Jester [^9] https://eigentaste.berkeley.edu/dataset/: The Jester dataset focuses exclusively on jokes. Users of the Jester online platform rate jokes and these ratings are then used to personalize joke recommendations for them.
  • Netflix [^10] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data: Netflix provided a training dataset consisting of 100,480,507 ratings from 480,189 users for 17,770 films. Each rating is represented as a set of four elements: ¡user, movie, rating date, rating score¿. The user and the movie are identified by integer IDs, and the rating scores range from 1 to 5 stars, also as integers.
  • Yahoo Music [^11] https://webscope.sandbox.yahoo.com/catalog.php?datatype=r: The Yahoo Music dataset is known for its large scale and diversity. It contains a large collection of user ratings on different musical elements such as tracks, albums, artists and genres. This dataset was used in the KDD-Cup 2011 competition, where participants were asked to analyse user preferences in music based on these ratings.
@@
  • Adressa [^14] https://reclab.idi.ntnu.no/dataset/ [178]: The Adressa dataset is a corpus of Norwegian news articles related to anonymous users. It is a collaborative project between the Norwegian University of Science and Technology and Adressavisen. The objective is to gain insight into the nature of news articles and their readers.
@@
  • Steam [^19] https://github.com/kang205/SASRec [183]: The Steam dataset is a collection of information about games published on the Steam platform. It includes details such as game names, release dates, genres, developers, publishers, and other relevant information.
  • TaFeng [^20] https://www.kaggle.com/datasets/chiranjivdas09/ta-feng-grocery-dataset: The TaFeng dataset is a comprehensive collection of supermarket shopping data, including detailed transaction records from the Ta Feng supermarket in Taiwan, covering a period from November 2000 to February 2001. The dataset comprises a variety of data points, including customer demographics, product categories, and detailed item descriptions along with quantities purchased.
@@
Both types of feedback play a critical role in the development of FRSs, providing diverse insights into user preferences that help improve the accuracy and relevance of the recommendations provided.
@@ VII-B Metrics
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions [185], such as Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio (HR), F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall.
Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision (AP) and Mean Average Precision (MAP) are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank (MRR) [186], Normalized Mutual Rank (NMR) [187] and Normalised Discounted Cumulative Gain (NDCG) [188] also contribute by assessing how well the top recommended items are ranked.
@@
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA) [189], which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate (CTR) [190] are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate (CVR) [190] measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index [191] evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
@@ VIII Conclusion
This study comprehensively examines the integration of FRSs with FMs, a direction that has gained attention for its ability to protect user privacy. The article begins by summarizing common approaches of FRSs and FMs, then delves into the challenges faced during integration. To address these challenges, the paper proposes various strategies including using transfer learning and multitask learning techniques to adapt to data diversity, employing privacy-preserving methods like differential privacy and secure multi-party computation, and reducing communication overhead through model compression and efficient communication protocols.
This work also discusses the future research directions directions indicating that FRSs can provide more accurate personalized recommendations while better protecting user privacy. Additionally, the paper showcases applications of FRS in various fields, demonstrating their potential and value in the real world. Through this study, we aim to provide theoretical guidance for integrating FRS with FM, directing future research and technological advancements to collectively advance this field.

### v3

@@ VI Future Directions
The emerging field of applying FMs to FRSs presents both opportunities and challenges, shown in Fig. 8. By leveraging the powerful representation capabilities of FMs and the privacy-preserving advantages of FRSs, it is possible to significantly enhance recommendation performance while addressing some of the long-standing challenges faced by traditional methods. This section discusses several key future directions for integrating FMs into FRSs, focusing on specific problems, FMs’ advantages, the new challenges introduced, and potential research directions. Sec. VI aims to guide future research and applications towards delivering higher-quality recommendation services while protecting user privacy.
@@ VI-A Data Augmentation
The challenge of data scarcity is a common problem in FRS due to data privacy constraints. Generative FMs, such as Large Language Models (LLMs) and Diffusion Models (DMs), can be utilized to alleviate this issue by generating synthetic user interaction data for training purposes, which can help mitigate data imbalance and scarcity [142, 143]. FM can generate virtual interaction records or additional item descriptions, thereby enhancing user profiles and improving recommendation accuracy and personalization [144].
However, the use of generated data introduces challenges related to data quality and potential biases, which may degrade the performance of recommendation systems. Rigorous validation and cleaning are necessary to ensure that the synthetic data is statistically consistent with real data, thereby reducing noise and biases [145]. Furthermore, integrating privacy-preserving techniques, such as differential privacy, can ensure that synthetic data generation does not reveal sensitive user information. Research should also focus on generating diverse data to cover a wide range of scenarios and user behaviors, thus reducing the risks of bias and overfitting [146].
@@ VI-B Cold-Start Recommendation
The cold-start problem, caused by a lack of sufficient historical interaction data for new users or items, presents a significant challenge for FRSs [147]. FMs, pre-trained on large amounts of textual data, possesses rich semantic knowledge that can be leveraged for generating high-quality representations of users and items, enabling zero-shot and few-shot learning to address cold-start scenarios [148, 38].
Despite FMs’ capabilities, ensuring user privacy in a federated environment remains challenging. Additionally, multi-task learning techniques can be explored to improve model adaptability and generalization across different domains, allowing the model to provide effective recommendations for new users and items while maintaining privacy [149]. Future research should focus on enhancing privacy-preserving mechanisms and exploring efficient knowledge transfer techniques to support cold-start recommendations.
@@ VI-C Multi-Modal Recommendation
The ability to process and integrate different modalities of data, such as text, images, audio, and video, is one of FM’s key strengths [10]. By incorporating multi-modal data, FM can construct more comprehensive user profiles [150], which in turn enhances the personalization of recommendation services [151, 152]. This richer profiling allows for more accurate user interest representation and better recommendation quality.
However, the heterogeneity of multi-modal data poses a challenge, as different data modalities have distinct structures and characteristics [153]. Developing unified representation methods to map different modalities into a common space is crucial for seamless model integration [150]. Privacy is also a concern, as multi-modal data often contains more sensitive information [154, 155]. Future research should focus on developing robust privacy-preserving techniques tailored for multi-modal data and improving computational efficiency in processing and integrating these diverse data types.
@@ VI-D Real-Time Recommendations
Real-time recommendation is critical in FRS, as it allows for dynamically adapting recommendations based on real-time user behavior and contextual data [156]. FM’s contextual understanding capabilities enhance the accuracy and relevance of real-time recommendations by effectively processing user queries and item descriptions.
However, real-time recommendations require extremely low latency, which places a high demand on computational resources. FM’s high computational requirements may introduce latency issues, adversely affecting the user experience. Future research should focus on model compression and acceleration techniques, such as distillation [157] and pruning [158], to reduce computational complexity and improve response times. Additionally, methods like sliding windows can be used to manage the input context efficiently, optimizing the handling of long sequences of user behavior.
VI-E Enhanced Explainability
Explainability is an essential aspect of FRS, as providing transparent explanations for recommendations increases user trust and satisfaction. Language FMs, pre-trained on extensive text data, are capable of generating coherent and informative explanations that can help users understand the reasons behind recommendations [159, 160]. For example, models such as ChatGPT can generate natural language explanations for why specific items are recommended [161].
The challenge, however, lies in the computational cost associated with generating these detailed explanations, especially for large language models, which can lead to latency issues [10]. Moreover, biases present in pre-training data may lead to unfair or biased explanations [28]. Future research should focus on model compression to reduce computational costs and on debiasing methods to minimize biases in generated explanations. Incorporating user feedback may further help improve the quality and fairness of recommendation explanations.
VI-F Advanced Metrics
The evaluation of FRSs with FMs requires advanced metrics that extend beyond traditional metrics like rating prediction or item ranking [162]. Given FM’s ability to generate new items and provide explanations, traditional metrics may not fully capture its performance. New evaluation metrics that account for the generative and explanatory capabilities of FMs are needed to better assess recommendation quality in these emerging scenarios [35].
Future research should explore comprehensive evaluation criteria that consider aspects such as diversity, fairness, and user satisfaction to provide a holistic assessment of FRS performance with FM integration.
@@ VII Resources
In this section, we introduce the commonly used datasets and evaluation metrics in FRS. The aim is to provide a comprehensive overview of the resources used in the field, helping readers understand the types of materials and benchmarks typically employed. This section presents various datasets, including explicit and implicit feedback types, along with metrics for evaluating recommendation performance. These datasets and metrics are crucial for assessing and improving the effectiveness of FRSs.
@@ VII-A Datasets
As shown in Fig. 9, the datasets used can be categorized into two types based on user feedback in FRS: explicit feedback and implicit feedback. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table I:
@@
  • Anime [^4] https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database: It collects user preference data from the MyAnimeList website, and contains information from 73,516 users on 12,294 anime titles. Users can add anime to their completed list and rate them, and this dataset compiles these ratings.
@@
  • Douban [^6] https://www.kaggle.com/datasets/utmhikari/doubanmovieshortcomments: The Douban Movie dataset is a Chinese website where internet users can post their opinions and comments about films. It contains over 2 million short comments on 28 movies from the Douban Movie website.
@@
  • Jester [^9] https://eigentaste.berkeley.edu/dataset/: It focuses exclusively on jokes. Users of the Jester online platform rate jokes and these ratings are then used to personalize joke recommendations for them.
  • Netflix [^10] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data: Netflix provided a training dataset consisting of 100,480,507 ratings from 480,189 users for 17,770 films. The user and the movie are identified by integer IDs, and the rating scores range from 1 to 5 stars, also as integers.
  • Yahoo Music [^11] https://webscope.sandbox.yahoo.com/catalog.php?datatype=r: It is known for its diversity and large scale. It contains a large collection of user ratings on different musical elements such as artists and genres.
@@
  • Adressa [^14] https://reclab.idi.ntnu.no/dataset/ [171]: It is a corpus of Norwegian news articles related to anonymous users. It is a collaborative project between the Norwegian University of Science and Technology and Adressavisen. The objective is to gain insight into the nature of news articles and their readers.
@@
  • Steam [^19] https://github.com/kang205/SASRec [176]: Steam is a collection of information about games published on the Steam platform. It includes details such as game names, release dates, genres, developers, publishers, and other relevant information.
  • TaFeng [^20] https://www.kaggle.com/datasets/chiranjivdas09/ta-feng-grocery-dataset: TaFeng is a comprehensive collection of supermarket shopping data, including detailed transaction records from the Ta Feng supermarket in Taiwan, covering a period from November 2000 to February 2001. The dataset comprises a variety of data points, including customer demographics, product categories, and detailed item descriptions along with quantities purchased.
@@
Both types of feedback play a critical role in the development of FRSs, providing diverse insights into user preferences that help improve the recommendation accuracy and relevance.
Figure 10: A potential unified framework for FRSs integrates server-side global aggregation and hyperparameter optimization with client-side data processing, model training, and evaluation.
@@ VII-B Metrics
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions [178], such as Mean Absolute Error, Mean Squared Error and Root Mean Squared Error. When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio, F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall. Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision and Mean Average Precision are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank [179], Normalized Mutual Rank [180] and Normalised Discounted Cumulative Gain [181] also contribute by assessing how well the top recommended items are ranked.
@@
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA) [182], which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate [183] are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate [183] measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index [184] evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
VII-C Open-source framework
Although there are already mature open-source frameworks for both RSs [185, 186, 187, 188] and FL [189, 190, 191], there remains a significant gap when it comes to FRSs. Currently, there is no mature, unified, and widely applicable open-source framework specifically designed for FRSs. Efforts such as FedMR [^22] https://github.com/mtics/FedMR [150] attempt to bridge this gap by integrating features from existing frameworks like MMRec [^23] https://github.com/enoche/MMRec [185] and RecBole [^24] https://github.com/RUCAIBox/RecBole [185], extending them into the realm of federated multimodal recommendation. However, FedMR is still in its infancy and does not yet provide the robustness or general applicability required of a comprehensive solution for FRSs.
As illustrated in Fig. 10, we make a preliminary attempt to propose a potential unified framework for FRSs in this paper, which aims to ideally encompass the following:
  • Integration of RSs and FL: Such a framework should seamlessly integrate the strengths of traditional recommendation systems with privacy-preserving capabilities from federated learning, addressing both data heterogeneity and privacy constraints effectively.
  • Modular and Extensible Architecture: The framework should feature a modular design that separates core components such as data preprocessing, model training, evaluation, and aggregation, facilitating improvements and the easy integration of new modules or functionalities.
  • Support for Multimodal Data: Leveraging the strengths of multimodal models, the framework should support the incorporation of various data modalities, enabling a more comprehensive understanding of user preferences and improving recommendation accuracy.
  • Scalability and Device Adaptability: The framework must be capable of scaling to accommodate a large number of clients and adapt to heterogeneous devices with varying computational capabilities, ensuring efficient performance on resource-constrained devices.
  • Privacy and Security: The framework should incorporate advanced privacy-preserving mechanisms, such as differential privacy and secure multi-party computation, to guarantee data confidentiality during both training and communication phases.
Despite the absence of a mature open-source FRS framework, the existing building blocks in RS and FL frameworks provide a strong foundation. By extending these capabilities and overcoming the unique challenges posed by FRSs, our proposed framework serves as a step towards a widely adopted, unified solution that fully supports FRSs.
@@ VIII Conclusion
This paper comprehensively examines the integration of FRSs with FMs, a direction that has gained attention for its ability to protect user privacy. It begins by summarizing common approaches of FRSs and FMs, then delves into the challenges faced during integration. To address these challenges, we comprehensively various strategies to adapt to data diversity, employing privacy-preserving methods, and reducing communication overhead through model compression and efficient communication protocols.
This work also discusses the future research directions directions indicating that FRSs can provide more accurate personalized recommendations while better protecting user privacy. Additionally, we showcase applications of FRS in various fields, demonstrating their potential and value in the real world. Through this paper, we aim to provide a guidance for integrating FRS with FM, directing future research and technological advancements to collectively advance this field.